# ReflecSched: Solving Dynamic Flexible Job-Shop Scheduling via LLM-Powered Hierarchical Reflection

### v1

@@ 4.3 The Pitfall of Myopic Greed
To quantify this behavior, we evaluated baseline LLMs using the Greedy Decision Ratio (GDR) on GEN-Bench, a benchmark specifically designed to control for structural biases that could otherwise mask or mimic such tendencies. The quantitative results (Figure 5) reveal that the GDR is persistently high across all tested models, indicating a significant greedy bias (Baeumel, van Genabith, and Ostermann 2025). A qualitative case study (Figure 1) further illustrates the consequences, showing how a single myopic choice can trigger a cascade of downstream inefficiencies. This combination of quantitative and qualitative data demonstrates that myopic behavior, stemming from the models’ autoregressive nature, is a critical pitfall for direct generation approaches in scheduling.
5 The ReflecSched Framework
@@ 5.1 Framework Overview
To address the identified limitations, we introduce ReflecSched, a novel framework that structurally decouples long-horizon, strategic reasoning from low-latency, online decision-making via two interconnected, sequential modules (Figure 4) (Yao et al. 2023b). The first, the Hierarchical Reflection Module, serves as the framework’s strategic planning component. It leverages multi-level, heuristic-driven simulations to explore future state trajectories (Yao et al. 2023a). It then employs the LLM to analyze these simulated outcomes, distilling them into a concise, actionable Strategic Experience $\mathcal{E}$. This experience then informs the second component, the Experience-Guided Decision-Making Module, which handles real-time execution. Guided by the strategic foresight in $\mathcal{E}$, this module makes a strategically-informed decision based on the immediate environment state.
This two-stage architecture is designed to address each of the identified challenges: First, by explicitly simulating future consequences, the Hierarchical Reflection Module provides the necessary foresight to counteract myopic greed, shifting the model’s focus from locally optimal choices to globally efficient strategies. Second, the framework directly addresses the underutilization of heuristics by employing PDRs to guide its exploratory rollouts (Zhang et al. 2024). This ensures that expert procedural knowledge is integrated into the reasoning process. Finally, the framework mitigates the long-context paradox. By having the Reflection Module distill its findings into a concise Strategic Experience $\mathcal{E}$, the final decision prompt supplied to the LLM contains only this high-level guidance and the immediate state (Jiang et al. 2024). This precludes informational overload and concentrates the model’s reasoning on the most salient data.
@@ 5.2 Hierarchical Reflection
Central to ReflecSched is the Hierarchical Reflection module, a mechanism designed to perform strategic evaluation, which is essential for the DFJSP. It implements a recursive “Simulate-Reflect-Refine” loop (Gui et al. 2025). This iterative process first explores the future through simulation, then reflects on the outcomes to identify key strategic principles, and finally refines these principles into a concise, actionable policy directive (Shinn et al. 2023).
@@
Finally, in the refinement phase, these two contrasting trajectories are provided to the LLM, which acts as a higher-level policy synthesizer (Light et al. 2025). Its task is not merely to summarize but to identify the key distinguishing features between $\tau^{(l)}_{best}$ and $\tau^{(l)}_{worst}$ and distill these observations into a concise, textual strategic guideline, termed the Strategic Experience $\mathcal{E}$. This synthesis, represented by the function $F_{LLM}$, effectively transforms high-dimensional, numerical simulation data into low-dimensional, human-interpretable strategic guidance:
@@ 5.3 Experience-Guided Decision-Making
Theoretically, this process connects to the principles of Approximate Policy Iteration (API) (Gordon 1995). The validity of this connection, and the resulting performance guarantee, hinges on two key working assumptions regarding the reflection process: (1) the rollout cost $\hat{J}$ serves as an approximation of the true cost-to-go function for the base policy $\pi_{\mathrm{base}}$; and (2) the LLM operates as a faithful synthesizer, meaning its generated experience $\mathcal{E}$ correctly reflects the superiority of the best simulated trajectory over the worst (Faithful Reflection).
The experience $\mathcal{E}_{t}$ generated under these conditions is then used to define the final policy $\pi_{\mathcal{E}_{t}}$, which makes a greedy decision with respect to the guidance (Schmied et al. 2025):
@@
With the policy thus defined, our framework inherits the classical policy-improvement guarantee of rollout algorithms (Bertsekas 2012):
Theorem 1 (Monotonic Improvement Guarantee).
Let $\pi_{\mathcal{E}_{t}}$ be the policy defined above. If the assumptions of cost function approximation and Faithful Reflection hold, then for all states $S_{t}$, its expected performance is no worse than that of the base heuristic policy, $\pi_{\mathrm{base}}$:
$V^{\pi_{\mathcal{E}_{t}}}(S_{t})\;\leq\;V^{\pi_{\mathrm{base}}}(S_{t})$ (6)
Figure 6: Comparative performance of ReflecSched and the LLM-Direct baseline on the GEN-Bench dataset. Bars (left y-axis) represent the average Relative Percent Deviation (RPD), where lower is better. The green line (right y-axis) shows the Win Rate of ReflecSched against the baseline. For reference, the performance of the single best-performing PDR (i.e., the heuristic with the lowest average RPD across all instances) and the average of all PDRs are shown as red and purple dashed lines, respectively. P-values from a Wilcoxon signed-rank test are annotated above each RPD bar group.
Theorem 1 establishes a conditional theoretical performance guarantee, grounding ReflecSched in established reinforcement learning principles and ensuring that, under plausible assumptions, its performance will not degrade below that of the underlying heuristics. While a formal proof of the underlying assumptions is intractable for black-box LLMs, our empirical study Section 6 shows a consistent and significant reduction in makespan. This suggests that these assumptions hold sufficiently well in practice for performance improvements to be realized. This combination of a conditional theoretical guarantee and strong empirical performance suggests that ReflecSched is an effective and well-grounded framework for dynamic scheduling.
@@ 6 Experiments
We conduct a series of experiments to evaluate the ReflecSched framework. In Section 6.1, we detail the experimental setup, which includes the benchmark datasets, evaluation metrics, hyperparameter tuning process, and the specific protocols used for different analyses. In Section 6.2, we present a comparative performance analysis to assess ReflecSched’s effectiveness against both the LLM-Direct baseline and traditional heuristics, and to empirically verify its mitigation of the identified pitfalls. In Section 6.3, we conduct an ablation study to isolate and quantify the contribution of the framework’s core hierarchical reflection mechanism.
@@ 6.1 Experimental Setup
Benchmark Datasets.
Our evaluation is performed on two benchmark suites we developed for this study: GEN-Bench, designed for general-purpose evaluation, and PDR-Bench, a diagnostic benchmark specifically constructed to assess heuristic utilization. Both suites contain instances at two scales: Normal (20 instances for GEN-Bench, 112 for PDR-Bench) and Small (20 for GEN-Bench, 12 for PDR-Bench). The detailed generation and curation process for these datasets is described in Appendix E.
Metrics and Models.
Our primary metrics are: (1) Relative Percent Deviation (RPD) (Huang, Gao, and Li 2025) calculated against the best makespan found for each instance across all methods and all runs evaluated in this study and (2) Win Rate, the percentage of instances where ReflecSched achieves a strictly better makespan than LLM-Direct. To assess the statistical significance of the performance difference, we use the non-parametric Wilcoxon signed-rank test (Demšar 2006). We evaluate a range of LLMs, including GPT-4o (Hurst et al. 2024), DeepSeek-V3 (Liu et al. 2024), and the Qwen3 series (8B, 14B, 32B) (Yang et al. 2025). All models are prompted in a zero-shot manner, utilizing Chain-of-Thought reasoning (Wei et al. 2022). The specific prompt structures, which encode shop-floor state and strategic guidance into natural language, are detailed in Appendix F.
Hyperparameter Tuning.
The core hyperparameters for ReflecSched are its reflection depth, Level ($L$), and its exploration breadth, the Number of Rollouts ($R$). To determine a suitable configuration, we conducted a systematic analysis on a dedicated validation set of 40 instances using the Qwen3-8B model, with average RPD as the performance metric. Based on this analysis, we selected the configuration of $L=6$ and $R=24$, which yielded the best average RPD on the validation set, for all subsequent experiments. Detailed setup and results of the hyperparameter analysis are provided in Appendix G.
Evaluation Protocol.
Our evaluation protocol is adapted to the specific goal of each experiment. For the initial diagnostic analysis (Sections 4.1-4.3), our aim is to capture the baseline model’s intrinsic behavioral tendencies. To achieve this, we employ a multi-sample voting scheme: actions are decided by the majority vote from 5 independent generations at a higher temperature ($T=0.8$) (Wang et al. 2022). In contrast, for the final comparative experiments designed to assess practical performance, all LLM inferences within the ReflecSched framework and for the LLM-Direct baseline use a low temperature ($T=0.2$) for efficient and stable decision-making. To mitigate outcome variability due to model stochasticity and provide a more stable performance estimate, we report the average results across 3 independent runs.
6.2 Comparative Performance Analysis
Our experimental goal is to primarily investigate the effectiveness of our hierarchical reflection architecture. Therefore, our main comparison is against the LLM-Direct baseline, with strong heuristics included as a classical performance reference. A direct comparison to DRL methods is omitted due to fundamental differences in methodology, as our work is motivated by circumventing their complex state engineering requirements.
@@ 6.2 Comparative Performance Analysis
ReflecSched demonstrates a statistically significant performance improvement over the LLM-Direct baseline across all tested configurations (Wilcoxon signed-rank test, $p<0.05$), securing a 71.35% average Win Rate and an average RPD reduction of 2.755% (see Figure 6 for details). Furthermore, the framework demonstrates the capability to outperform all evaluated PDR, with several of the tested LLM backbones enabling performance superior to any individual heuristic. For a further comparison against a baseline constructed by selecting the best-performing heuristic for each instance, the resulting average RPDs are minimal, and a two-sided Wilcoxon signed-rank test confirms no statistically significant difference between their performances (Table 1, $p>0.05$).
The framework’s superior performance can be explained by its success in mitigating the specific pitfalls identified in our analysis. First, it addresses the underutilization of heuristics, with results on the diagnostic PDR-Bench (Figure 3) showing that ReflecSched consistently and substantially reduces the RPD compared to the baseline across all evaluated models. Second, it directly counteracts myopic greed; the data in Figure 5 shows a consistent reduction in the GDR across all models, providing strong evidence that the generated strategic experience steers the LLM towards non-myopic decisions. Finally, it circumvents the long-context paradox. By having the reflection module distill complex state information into a concise experience $\mathcal{E}$, the final decision prompt becomes substantially shorter and more focused. This is empirically validated by our token consumption analysis (Appendix H), which shows that on the larger, more complex Normal-scale instances, ReflecSched requires fewer tokens than the baseline. This efficiency gain is less pronounced on simpler Small-scale problems, where the benefits of such strategic distillation offer naturally diminished returns.

### v2

@@ 4.3 The Pitfall of Myopic Greed
To quantify this myopic behavior, we introduce the Greedy Decision Ratio (GDR). We formally define the one-step greedy action at any decision point $t$ as the one that minimizes the local finish time:
$a^{\text{greedy}}_{t}=\arg\min_{a\in\mathcal{A}(S_{t})}(\text{est}(a)+p(a))$ (15)
where $\text{est}(a)$ is the earliest start time of an action and $p(a)$ is its processing time. The GDR is then the fraction of decision points where the agent’s chosen action matches $a^{\text{greedy}}_{t}$.
We evaluated baseline LLMs using this metric on GEN-Bench, a benchmark specifically designed to be globally balanced and control for what we term rule-balance bias—a scenario where a dataset might implicitly favor a single, simple heuristic. This curation is critical: on a rule-balanced benchmark, a persistently high GDR (e.g., $\approx$80-90%) is a strong indicator of intrinsic myopic decision-making rather than an artifact of the model correctly identifying a simple optimal heuristic for the dataset. The quantitative results (Figure 4) reveal that the GDR is indeed persistently high across all tested models, indicating a significant greedy bias [baeumel2025lookaheadlimitationmultioperandaddition]. A qualitative case study (Figure 1) further illustrates the consequences, showing how a single myopic choice can trigger a cascade of downstream inefficiencies. Specifically, the top schedule makes a locally optimal choice by assigning operation $O_{1,1}$ to Machine $M_{3}$ for an earlier finish time, but this leads to a less balanced machine workload and a downstream bottleneck, ultimately resulting in a suboptimal final makespan (7.51 vs. 6.51). This combination of quantitative and qualitative data demonstrates that myopic behavior, stemming from the models’ autoregressive nature, is a critical pitfall for direct generation approaches in scheduling.
@@
5 ReflecSched Framework
@@ 5.1 Framework Overview
To address the identified limitations, we introduce ReflecSched, a novel framework that structurally decouples long-horizon, strategic reasoning from low-latency, online decision-making via two interconnected, sequential modules (Figure 5[yao2023react]. The first, the Hierarchical Reflection Module, serves as the framework’s strategic planning component. It leverages multi-level, heuristic-driven simulations to explore future state trajectories [yao2023tree]. It then employs the LLM to analyze these simulated outcomes, distilling them into a concise, actionable Strategic Experience $\mathcal{E}$. This experience then informs the second component, the Experience-Guided Decision-Making Module, which handles real-time execution. Guided by the strategic foresight in $\mathcal{E}$, this module makes a strategically-informed decision based on the immediate environment state.
This two-stage architecture is designed to address each of the identified challenges: First, by explicitly simulating future consequences, the Hierarchical Reflection Module provides the necessary foresight to counteract myopic greed, shifting the model’s focus from locally optimal choices to globally efficient strategies. Second, the framework directly addresses the underutilization of heuristics by employing PDRs to guide its exploratory rollouts [zhang2024restmctsllmselftrainingprocess]. This ensures that expert procedural knowledge is integrated into the reasoning process. Finally, the framework mitigates the long-context paradox. By having the Reflection Module distill its findings into a concise Strategic Experience $\mathcal{E}$, the final decision prompt supplied to the LLM contains only this high-level guidance and the immediate state [jiang2024longllmlinguaacceleratingenhancingllms]. This precludes informational overload and concentrates the model’s reasoning on the most salient data.
@@ 5.2 Hierarchical Reflection
Central to ReflecSched is the Hierarchical Reflection module, a mechanism designed to perform strategic evaluation, which is essential for the DFJSP. It implements a recursive “Simulate-Reflect-Refine” loop [gui2025hypertreeplanningenhancingllm]. This iterative process first explores the future through simulation, then reflects on the outcomes to identify key strategic principles, and finally refines these principles into a concise, actionable policy directive [shinn2023reflexion].
@@
Finally, in the refinement phase, these two contrasting trajectories are provided to the LLM, which acts as a higher-level policy synthesizer [light2025strategistselfimprovementllmdecision]. Its task is not merely to summarize but to identify the key distinguishing features between $\tau^{(l)}_{best}$ and $\tau^{(l)}_{worst}$ and distill these observations into a concise, textual strategic guideline, termed the Strategic Experience $\mathcal{E}$. This synthesis, represented by the function $F_{LLM}$, effectively transforms high-dimensional, numerical simulation data into low-dimensional, human-interpretable strategic guidance:
@@ 5.3 Experience-Guided Decision-Making
Theoretically, this process connects to the principles of Approximate Policy Iteration (API) [Gordon1995StableFunction]. The validity of this connection, and the resulting performance proposition, hinges on two key working assumptions regarding the reflection process: (1) the rollout cost $\hat{J}$ serves as a reasonable approximation of the true cost-to-go function for the base policy $\pi_{\mathrm{base}}$; and (2) the LLM operates as a faithful synthesizer, meaning its generated experience $\mathcal{E}$ correctly reflects the superiority of the best simulated trajectory over the worst (Faithful Reflection).
The experience $\mathcal{E}_{t}$ generated under these conditions is then used to define the final policy $\pi_{\mathcal{E}_{t}}$, which makes a greedy decision with respect to the guidance [schmied2025llmsgreedyagentseffects]:
@@
With the policy thus defined, our framework can be understood through the lens of classical rollout algorithms [bertsekas2012dynamic], suggesting a conditional improvement over the base policy.
Proposition 1 (Conditional Policy Improvement).
Let $\pi_{\mathcal{E}_{t}}$ be the policy defined above and $V^{\pi}(S_{t})$ be the expected makespan from state $S_{t}$ under policy $\pi$. If the assumptions of Cost Function Approximation and Faithful Reflection hold, then for all states $S_{t}$, the expected performance of $\pi_{\mathcal{E}_{t}}$ is no worse than that of the base heuristic policy, $\pi_{\mathrm{base}}$. This relationship holds over the expectation of stochastic events in the environment and the randomness of the base policy:
$\mathbb{E}[V^{\pi_{\mathcal{E}_{t}}}(S_{t})]\;\leq\;\mathbb{E}[V^{\pi_{\mathrm{base}}}(S_{t})]$ (21)
Proposition 1 establishes a conditional theoretical proposition, grounding ReflecSched in established reinforcement learning principles. It suggests that, under plausible assumptions, the framework’s performance is not expected to degrade below that of the underlying heuristics. It is important to note that the simulated lookahead assumes no new dynamic events occur within its finite planning horizon. While a formal proof of the underlying assumptions is intractable for black-box LLMs, our empirical study in Section 6 shows a consistent and significant reduction in makespan. This suggests that these assumptions hold sufficiently well in practice for performance improvements to be realized. This combination of a conditional theoretical foundation and strong empirical performance suggests that ReflecSched is an effective and well-grounded framework for dynamic scheduling.
@@ 6 Experiments
Figure 6: Comparative performance of ReflecSched and the LLM-Direct baseline on the GEN-Bench dataset. Bars (left y-axis) represent the average RPD, where lower is better. The green line (right y-axis) shows the Win Rate of ReflecSched against the baseline, indicating the percentage of instances where it finds a strictly better solution. For reference, the performance of the single best-performing PDR (i.e., the heuristic with the lowest average RPD across all instances) and the average of all PDRs are shown as red and purple dashed lines, respectively. Asterisks denote statistical significance from Wilcoxon signed-rank tests (* p ¡ 0.05, ** p ¡ 0.01, *** p ¡ 0.001).
We conduct a series of experiments to rigorously evaluate our proposed ReflecSched framework. In Section 6.1, we detail the experimental setup, including the benchmark datasets, evaluation metrics, and models. In Section 6.2, we present a comparative performance analysis of ReflecSched against baseline methods and traditional heuristics. In Section 6.3, we conduct ablation studies to isolate the contribution of the framework’s core components and analyze the sensitivity of its key hyperparameters. Finally, in Section 6.4, we analyze the computational cost of the framework in terms of token consumption.
@@ 6.1 Experimental Setup
Benchmark Datasets. Our evaluation is performed on two benchmark suites developed for this study: GEN-Bench, for general-purpose evaluation, and PDR-Bench, a diagnostic benchmark to assess heuristic utilization. Both suites contain instances at two scales: Normal and Small. The detailed generation and curation methodology for these datasets is described in D.
Metrics and Models. To ensure a fair and scalable comparison across instances of varying difficulty, we use relative performance metrics. Our primary metrics are:
  1. 1.
    Relative Percent Deviation (RPD), which normalizes performance against the best-known solution for each instance, calculated as:
    $\text{RPD}(\%)=\frac{C_{\max}^{\text{method}}-C_{\max}^{\text{best}}}{C_{\max}^{\text{best}}}\times 100$ (22)
    where $C_{\max}^{\text{method}}$ is the makespan from the evaluated method and $C_{\max}^{\text{best}}$ is the best makespan found for that instance across all methods and runs in this study. Lower RPD is better.
  2. 2.
    Average Rank, which measures the consistency of a method’s performance. For each instance, all competing methods are ranked based on their makespan. A method’s average rank is its mean rank across all instances. A lower average rank indicates a consistently superior method.
To assess statistical significance, we use the non-parametric Wilcoxon signed-rank test [Janez2006Statistical]. We evaluate a range of LLMs, including GPT-4o [openai2024gpt4ocard], DeepSeek-V3 [deepseekai2025deepseekv3technicalreport], and the Qwen3 series (8B, 14B, 32B) [yang2025qwen3technicalreport]. All models are prompted in a zero-shot manner with Chain-of-Thought reasoning [wei2022chain].
Evaluation Protocol. For initial diagnostic analyses aimed at capturing intrinsic model behaviors, we employ a multi-sample voting scheme: actions are decided by a majority vote from 5 independent generations at a high temperature (T=0.8T=0.8[wang2022self]. For the final comparative experiments against the LLM-Direct baseline, all LLM inferences use a low temperature ($T=0.2$) for efficiency and stability. To mitigate stochasticity, we report the average results across 3 independent runs.
@@
Table 3: Comparison of our ReflecSched framework against baseline methods on Normal and Small datasets. Performance is evaluated using Average Relative Percentage Deviation (RPD, %) and Average Rank. Lower values are better for both metrics.
Category Method / Model Normal Dataset Small Dataset Overall Average
Avg RPD (%) Avg Rank Avg RPD (%) Avg Rank Avg RPD (%) Avg Rank
Baseline Methods NHS 8.27 5.70 13.42 6.10 10.85 5.90
CWH 7.84 5.35 13.86 4.75 10.85 5.05
DDQN-JS 13.10 6.40 4.91 3.30 9.01 4.85
ReflecSched (Ours) Qwen3-8b 3.38 3.50 10.01 4.15 6.69 3.83
Qwen3-14b 3.75 3.70 8.33 3.25 6.04 3.48
Qwen3-32b 3.95 3.75 8.29 2.60 6.12 3.18
GPT-4o 4.08 3.70 10.24 3.55 7.16 3.63
Deepseek-v3 4.86 3.75 9.89 3.35 7.38 3.55
6.2 Comparative Performance Analysis
6.2.1 Comparison with LLM-Direct and Heuristics
Our primary comparison is against the LLM-Direct baseline to investigate the effectiveness of our hierarchical reflection architecture. As shown in Figure 6, ReflecSched demonstrates a statistically significant performance improvement over LLM-Direct across all tested configurations. It achieves a 71.35% average Win Rate and an average RPD reduction of 2.755%.
Furthermore, the framework is capable of outperforming any single heuristic. We also compared its performance to an oracle baseline constructed by selecting the best-performing heuristic for each instance individually. As shown in Table 2, a two-sided Wilcoxon signed-rank test confirms no statistically significant difference between ReflecSched and this strong oracle baseline ($p>0.05$), highlighting the framework’s effectiveness.
6.2.2 Comparison with Baseline Methods
To benchmark our ReflecSched framework, we selected and adapted three advanced algorithms representing a spectrum of established paradigms. The novelty of our problem’s specific dynamic event types necessitated this adaptation, as no directly comparable open-source implementations exist. Our comparative analysis, therefore, is not intended to challenge the original results of these baselines in their native domains, but to rigorously demonstrate the practical advantages of our proposed LLM-driven approach in this specific, complex scheduling context.
The first, CWH (Composite Weighted Heuristic), represents a sophisticated heuristic approach. Unlike the single-criterion dispatching rules evaluated previously in our study like Table 2 and Figure 6, CWH calculates job priority by integrating multiple system features through a static, linear weighted sum, providing a strong non-learning baseline. Building on this, NHS (Neural Heuristic Scorer) advances the heuristic concept by replacing the static linear weights of CWH with a neural network, allowing it to learn a non-linear scoring function and adapt its policy based on training data. Moving into the domain of DRL, we included DDQN-JS, a single-agent DRL algorithm that learns an end-to-end scheduling policy directly from environmental interactions using a Double Deep Q-Network [liu2022deep].
The performance of ReflecSched, instantiated with five different LLMs, was compared against these baselines. The results, summarized in Table 3, unequivocally demonstrate the superior performance of our framework. Our best-performing variants, ReflecSched+Qwen3-14b and ReflecSched+Qwen3-32b, achieved the best overall average RPD and average rank, respectively. This represents a substantial performance leap over the strongest baseline, DDQN-JS, which only attained an overall average RPD of 9.01%. This underscores the significant advantage provided by our framework’s planning and reflection capabilities.
On the Normal Dataset, the performance gap is most pronounced. All ReflecSched instances achieve an RPD of 4.86% or lower, whereas the best-performing baseline on this dataset CWH could only reach 7.84%. The results on the Small Dataset reveal a more nuanced picture; while DDQN-JS is competitive in terms of RPD, our ReflecSched+Qwen3-32b variant achieves the absolute best average rank across all methods and datasets, highlighting its ability to generate consistently high-quality solutions.
In summary, the ReflecSched framework not only delivers superior overall performance but also demonstrates remarkable robustness. Unlike the baseline methods, whose effectiveness varies considerably between problem scales, ReflecSched consistently delivers top-tier results, proving its strong generalization and balanced performance.
6.3 Ablation and Sensitivity Analysis
All experiments in this section were conducted using Qwen3-8B as the base model.
6.3.1 Impact of Hierarchical Reflection
To isolate the contribution of our core architectural design, we conducted an ablation study comparing the full ReflecSched framework ($L=6,R=24$) against an ablated, single-level version ($L=0$). The results in Figure 7 confirm that the strategic foresight gained from multi-level reflection is a primary driver of the framework’s effectiveness. More revealingly, simply increasing the search breadth for the single-level model fails to improve its performance, which plateaus and can even degrade. This suggests that a brute-force, “flat” search is ineffective without hierarchical guidance and confirms that the hierarchical reflection mechanism provides a qualitative benefit that simply increasing computation cannot replicate.